Earlier reports Convert Day of Year

The Hubble Space Telescope

Unlocking the Secrets of the Universe

Daily Report #5350

Period Covered:
07:00 pm March 7, 2011 - 06:59 pm March 8, 2011
(DOY 67/0000z - 67/2359z)

Flight Operations Summary

Significant Spacecraft Anomalies: (The following are preliminary reports
of potential non-nominal performance that will be investigated.)

12747 - SCI: REAcq(2,1,1) at 067/02:50:36 and at 067/03:58:22 both fails to RGA Hold, search radius
limit exceeded on FGS-2@066/0252z
Observations affected: WFC3 24-28 Proposal ID#12379

12749 - SCI: GSAcq (2,0,2) failed to RGA Hold@067/1502z
Observation affected, Astrometry Proposal ID#12441

19037-0 - WFC3 RT Recovery Post-NSSC-1 9.2.0 Installation@67/1604z

19038-1 - Inhibit NSSC1 ATP and transition WFC3 to TECSAFE@67/1310z

19039-4 - Halt NSSC1, load BQ9.2 FSW and recover to Normal@67/1310z

19040-0 - Uplink N02Z1240C and set ATP pointer@67/1310z

19044-0 - Real Time OBAD (with maneuver)@67/0413z

19046-0 - Safe the ESM @67/1342z


FGS GSAcq		7		6
FGS REAcq		9		7
OBAD with Maneuver	7		7

The NSSC-1 BL/B! 9.2.0 FSW image was successfully transferred from the SSM486 via OR 19039-4 at 
067/14:36 UTC.  This was followed by a successful recovery of the NSSC-1 to Normal mode at 
067/14:40 UTC.    
The SIs have all been successfully recovered to their nominal Operate states following the 
installation of the NSSC-1 BL/BQ 9.2.0 FSW image earlier this morning (with the exception of 
NICMOS, which will continue to remain in Safe with TECs off).    
Real-time commanding to recover both WFC3 and the ESM, NCS CPL, & PCE was successfully completed at 
067/16:04 UTC and 067/17:10 UTC respectively.  Stored commanding recovery transitions from the SMS 
have been monitored since ~067/19:00 UTC through ~067/22:00 returning ACS, STIS, and COS, to their 
nominal boundary states. WFC3 will be the first instrument to return to nominal science operations 
with its first observation, scheduled for ~067/22:59 UTC.  The ACS WFC will follow at ~068/00:05 
UTC.  STIS preparations for return to science will continue through the LV turn-on of both MAMAs, 
which will complete tomorrow at ~068/15:00 UTC.  COS is continuing to thermally stabilize before 
resuming observations at ~068/22:39 UTC.   Because the ACS SBC has been off for several months, the 
068 SMS includes special commanding to perform a slow, full HV ramp-up (the final ramp-up from the 
recovery from anomalous shutdown procedure) starting at 069/12:44 UTC.  ACS Event Flag 2 in the 
NSSC-1 has already been cleared to allow this commanding to execute.  The ramp-up is followed two 
days later by an SBC Fold Test, with the first nominal ramp-up commanding starting at 071/20:29 UTC.

Program Principal Investigator Program Title
Science Observations Scheduled
12166 Harald Ebeling, University of Hawaii A Snapshot Survey of The Most Massive Clusters of Galaxies
12441 Edmund Nelan, Space Telescope Science Institute FGS Science During NSSC-1 Flight Software Upgrade
Calibration Observations Scheduled
12342 Tiffany Borders, Space Telescope Science Institute WFC3 UVIS CCD Daily Monitor
12344 Tiffany Borders, Space Telescope Science Institute Cycle 18: UVIS Bowtie Monitor
12379 Kai Noeske, Smithsonian Institution Astrophysical Observatory UVIS CTE Monitor: Star Clusters
12400 Michael Wolfe, Space Telescope Science Institute Dark Monitor Part 1
12402 Michael Wolfe, Space Telescope Science Institute CCD Bias and Read Noise Monitor-Part 1
Other Observations Scheduled
12046 David Sahnow, The Johns Hopkins University COS FUV DCE Memory Dump